# Feedback-Enhanced Online Multiple Testing with Applications to Conformal Selection

### v1

@@ 3.3 Optimized online conformal multiple testing with feedback
Under the same assumptions of Theorem 3, the Optimized Online Conformal Testing with Feedback by criterion ${\mathcal{M}}^{\rm EWMA}$ (Opt-OCTF) procedure in Algorithm 3 satisfies $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}.$
4 Numerical Simulations
@@
In this section, we present extensive synthetic experiments to demonstrate the validity and efficiency of our proposed methods. First, Section 4.1 reports results for online multiple testing in the non-conformal setting. Second, Sections 4.2 and 4.3 present the results under the conformal setting for OCTF and Opt-OCTF, respectively. For the conformal setting, we focus on binary classification scenarios; additional results, including a regression scenario, are provided in Appendix F. Following the setup in Robertson et al. [2023], we fix the pre-specified parameters as $\lambda=0.5$, $\gamma_{j}\propto j^{-1.6}$ for all $j\in\mathbb{N}$, and $s_{0}=\alpha/2$, and use these settings consistently across all experiments.
@@ 4.1 Synthetic experiments under non-conformal settings
We begin with experiments in traditional online multiple testing problems, covering three distinct scenarios:
  • Scenario I (Testing with Gaussian observations) We simulate $T$ independent test statistics $Z_{t}\sim N(\mu_{t},1)$ with hypotheses $\mathbb{H}_{0t}:\mu_{t}=0$. One-sided $p$-values are given by $p_{t}=\Phi(-Z_{t})$, where $\Phi$ is the standard Gaussian CDF. The signal strengths $\mu_{t}$ follow a mixture model:
@@
    where the random variable $F_{1}\sim N(\mu_{c},1)$ and $\mu_{c}=2.5$.
  • Scenario II (Testing with Beta alternatives) We generate $T$ $p$-values according to the following model:
@@
  • Scenario III (Testing under local dependence) We simulate $T$ correlated test statistics $(Z_{1},\dots,Z_{T})\sim N(\bm{\mu},\Sigma)$ where $\bm{\mu}=(\mu_{1},\ldots,\mu_{T})$ with $\mu_{t}=2$ for randomly a fraction $\pi_{1}$ of indices and $\mu_{t}=0$ otherwise. The covariance matrix $\Sigma$ has a block-diagonal structure: coordinates are grouped into blocks of size $n_{\mathrm{block}}=10$, with within-block correlation $\rho=0.8$ and diagonal elements as 1. We test hypotheses $\mathbb{H}_{0t}:\mu_{t}=0$ using one-sided $p$-values $p_{t}=\Phi(-Z_{t})$.
We set $T=1000$ and the FDR level at $\alpha=0.1$. We compare our methods SF and LF with state-of-the-art algorithms for online FDR control, namely LOND [Javanmard and Montanari, 2015], LORD++ [Ramdas et al., 2017], and SAFFRON [Ramdas et al., 2018], using their default parameters ($\lambda=0.5$ chosen for SAFFRON). In the locally dependence setting, we additionally include the proposed $\textbf{SF}_{\text{dep}}$ and $\textbf{LF}_{\text{dep}}$ along with the existing $\textbf{SAFFRON}_{\text{dep}}$ and $\textbf{LORD}_{\text{dep}}$ from Zrnic et al. [2021]. We evaluate the performance via empirical FDR and power averaged over 500 independent replications. Appendix F confirms that mFDR and FDR show similar trends.
@@
Figure 2 shows results for varying non-null proportion $\pi_{1}\in[0.1,0.8]$ under Scenarios I and II. All benchmark methods ensure valid FDR control across different $\pi_{1}$ under both Scenarios. Our feedback-based SF and LF methods enhance detection power while maintaining valid online FDR control: SF significantly outperforms SAFFRON, while LF yields higher power than LORD++ and surpasses SAFFRON when $\pi_{1}$ is small. In contrast, both LORD++ and LOND remain conservative across different values of $\pi_{1}$.
[…]
Figure 2: Results for Scenario I and Scenario II. Line charts of FDR and Power at stopping time with varying non-null proportion $\pi_{1}$ from $0.1$ to $0.8$ after $500$ replications; The black dashed lines denote the FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
@@
Figure 3: Results for Scenario III: FDR and power at stopping time across 500 replications with non-null proportion $\pi_{1}$ ranging from 0.1 to 0.8. The black dashed line indicates the target FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
@@ 4.2 Results for online conformal testing with feedback
  • Scenario IV: Data is generated as $\mathbf{X}\mid Y=0\sim\mathcal{N}_{4}\left(\bm{\mu}_{1},\mathbf{I}_{4}\right)$, and $\mathbf{X}\mid Y=1\sim\mathcal{N}_{4}\left(\bm{\mu}_{2},\mathbf{I}_{4}\right)$, where $\bm{\mu}_{1}=(1,0,0,0)^{\top},\bm{\mu}_{2}=(0,0,-2,-2)^{\top}$. The target region is $\mathcal{A}=\{1\}$.
We set the pre-specified stopping time to $T=1000$ and target $\operatorname{FDR}$ level to $\alpha=0.2$. The predictive model is a neural network trained on $n_{\text{tr}}=1000$, and the initial calibration set has size $n_{\text{cal}}=1000$. Appendix F presents additional results for varying $n_{\text{cal}}$ and alternative training algorithms $\widehat{\mu}$, along with the corresponding results for a regression task.
All methods are implemented within our proposed OCTF workflow; the only difference lies in how the sequence $\{\alpha_{t}\}$ is generated for each benchmark. For clarity, we refer to each method by its corresponding $\alpha_{t}$-generating algorithm and omit the “OCTF” prefix. We compare our proposed methods (SF, SFS, LF, and LFS) with existing approaches: SAFFRON, LORD++, and LOND. Performance is evaluated in terms of empirical FDR and power at $T$, averaged over 500 replications.
Figure 4 reports the online FDR and power at the stopping time $T$ under Scenario IV across varying non-null proportions $\pi_{1}\in[0.1,0.8]$. All methods control the FDR below the nominal level $\alpha$, with SF aligning most closely with the target level among all competitors. In terms of power, as expected, SF consistently achieves the highest power, while LF also performs competitively and attains higher power than SAFFRON across all $\pi_{1}$. The LOND algorithm exhibits the lowest power. Thanks to the incorporation of feedback—reflected in the term $\sum_{j\leq t}\gamma_{t-j}\alpha_{j}\theta_{j}$ within the testing levels—SFS and LFS outperform LORD++, despite discarding certain rejections under the non-null to ensure theoretical guarantees. This safe strategy mitigates the risk of FDR inflation in complex real-data scenarios while still leveraging feedback to enhance power.
Figure 4: Results for Scenario IV: the values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ at stopping time $T$ across different non-null proportion $\pi_{1}$. The black dashed lines denote the FDR level $\alpha=0.2$. Shaded areas show $\pm 1$ standard error.
@@ 4.3 Results for optimized online conformal testing with feedback
We now demonstrate the effectiveness of the proposed model optimization strategy from Section 3.3 for selecting predictive models in online conformal testing with FDR control, under scenarios where the non-null distribution shifts over time. Additional results for scenarios without distribution shifts are presented in Appendix F. We consider a binary classification example with the pre-specified stopping time $T=1000$.
  • Scenario V: Data are generated as $\mathbf{X}\mid Y=0\sim\mathcal{N}_{4}\left(\bm{\mu}_{1},\mathbf{I}_{4}\right)$, and $\mathbf{X}\mid Y=1\sim\mathcal{N}_{4}\left(\bm{\mu}_{2},\mathbf{I}_{4}\right)$, where $\bm{\mu}_{1}=(2,0,0,0)^{\top},\bm{\mu}_{2}=(0,0,-2,-2)^{\top}$. The target region is $\mathcal{A}=\{1\}$.
@@
For the historical labeled dataset, the non-null proportion is fixed as we fix $\pi_{t}=\operatorname{Pr}(Y_{t}=1)=0.2$. For the testing data, we consider a sine pattern shifts, where the non-null proportion varies as $\pi_{t}=\{\sin(8\pi t/T)+1\}/4$, oscillating between 0 and 0.5. We consider $K=3$ candidate models: random forest (RF), neural network (NN), and support vector machine (SVM). The model evaluation criterion is based on ${\mathcal{M}}^{\text{EWMA}}_{t}(k,{\mathcal{D}}_{t})$ with exponential weight parameter $\rho=0.95$. The target FDR level is $\alpha=0.1$. To reduce computational cost, we restrict the auxiliary non-null $p$-value computation to a sliding window of fixed length $L=200$: for $t>L$, we define ${\mathcal{C}}_{1t}=\{i:t-L\leq i<t,\theta_{i}=1\}$; for $t\leq L$, we use the full set ${\mathcal{C}}_{1t}=\{i<t:\theta_{i}=1\}$.
We evaluate several methods that, at each time $t$, select a conformity score from $K$ candidates, compute the corresponding conformal $p$-values, and apply a specific multiple testing procedure. In particular, we consider four proposed procedures (SF, SFS, LF, and LFS), each in two variants: an optimized version (“Opt-”) using our model selection strategy, and a random version (“Ran-") with randomly selected scores. As benchmarks, we include SAFFRON and LORD++, both using random score selection.
@@
The results for Scenario V are shown in Figure 5. All methods successfully control the empirical FDR around the nominal level $\alpha=0.1$. In terms of power, the optimized (Opt) variants consistently outperform their randomly selected (Ran) counterparts, while both Opt-LFS and Ran-LFS achieve higher power than LORD++. Both Opt-SFS and Ran-SFS procedures are more conservative, exhibiting lower power than SAFFRON due to discarding rejections under the non-null to ensure theoretical guarantees. Although this strategy reduces power in this synthetic setting, we will later observe that Opt-SF may suffer from FDR inflation on real data due to the lack of theoretical guarantees.
Overall, these results highlight the effectiveness of our model optimization strategy. Feedback-enhanced methods achieve significantly higher power than feedback-ignoring benchmarks. In contrast, random model selection results in a substantial loss of power, underscoring the importance of informed model choice particularly under distribution shifts.
@@
Figure 5: Results for Scenario V (sine pattern shifts): the values of $\operatorname{FDR}(t)$ and $\operatorname{Power}(t)$ across different time $t$. The black dashed lines denote the FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.

### v2

@@ 4.3 Optimized online conformal testing with feedback
Under the assumptions in Theorem 4, the Opt-OCTF procedure by ${\mathcal{M}}^{\rm EWMA}_{t}$ satisfies $\operatorname{mFDR}(t)\leq\alpha$ for all $t\in\mathbb{N}.$
Although Corollary 4.1 is stated for the EWMA criterion, the underlying framework for optimized online conformal testing can be extended more broadly. In particular, we allow more general selection criteria $\mathcal{M}$} by using a full permutation strategy over ${\mathcal{C}}_{0t}\cup\{t\}$; see Appendix C.2. In addition, data-driven score functions obtained by online updating are also compatible with the framework under a suitable symmetry-preserving training scheme. However, we focus on selection among $K$ pre-trained candidates in the main text for computational reasons; see Appendix C.3 for details.
Remark 4.1.
In (13), we include the current time point $t$ in the null calibration set to preserve symmetry over ${\mathcal{C}}_{0t}\cup\{t\}$, which is essential for the validity of the optimized conformal $p$-value after score selection. Although this may slightly affect calibration when $t$ is non-null, the effect is typically minor in practice. A simple truncated variant can further reduce this influence while preserving validity; see Appendix E.5 for details.
4.4 Optimality analysis of EWMA criterion
We now give a supporting analysis for the EWMA criterion by establishing certain optimality of the online score selection step. To facilitate analysis, we work under the standard two-group model, which is widely used to study the performance of online testing procedures [Gang et al., 2023, Humbert et al., 2025].
Assumption 2 (Two-group model).
Suppose the score functions follow the two-group model for each $k\in[K]$:
$V({\bf X}_{t};k)\overset{\rm indep}{\sim}\pi_{t}F_{k}^{(0)}+(1-\pi_{t})F^{(1)}_{k,t},\quad\theta_{t}\overset{\rm indep}{\sim}\text{Bernoulli}(1-\pi_{t}),$
where $F_{k}^{(0)}$ and $F^{(1)}_{k,t}$ are the null and (time-varying) non-null score distributions for $k$-th model.
Under this model, the oracle conformal $p$-value associated with score $V(\cdot;k)$ is $F_{k}^{(0)}(V({\bf X}_{t};k))$. Conditioned on $\theta_{t}=1$, smaller value of $F_{k}^{(0)}(V({\bf X}_{t};k))$ indicates higher detection power. And we define the optimal score function at time $t$ as
$k_{t}^{*}=\arg\min_{k\in[K]}{\mathbb{E}}[F_{k}^{(0)}(V({\bf X}_{t};k))\mid\theta_{t}=1].$
Note that ${\mathbb{E}}[F_{k}^{(0)}(V({\bf X}_{t};k))\mid\theta_{t}=1]$ is the population analogue of the empirical proxy ${\mathbb{E}}[p_{k,t}\mid\theta_{t}=1]$. To show that the EWMA criterion tracks this oracle target and selects ${k}^{*}_{t}$ reliably, we impose standard regularity conditions.
Assumption 3 (Slowly varying distribution).
For each $k\in[K]$ and $t\in\mathbb{N}$, $F^{(1)}_{k,t}$ has a slowly varying distribution $\|F^{(1)}_{k,t}-F^{(1)}_{k,t-1}\|_{\infty}\leq\gamma$ such that $\gamma\to 0$.
Assumption 4 (Non-degeneracy).
There exist constants $0<\underline{\pi}\leq\overline{\pi}<1$ such that null proportion satisfies $\pi_{t}\in[\underline{\pi},\overline{\pi}]$.
Assumption 5 (Margin).
For any $t$ and a constant $c>0$, $\min_{i\neq k_{t}^{*}}|{\mathbb{E}}[F_{i}^{(0)}(V({\bf X}_{t};i))\mid\theta_{t}=1]-{\mathbb{E}}[F_{k_{t}^{*}}^{(0)}(V({\bf X}_{t};k_{t}^{*}))\mid\theta_{t}=1]|\geq c$.
Assumption 3 allows the non-null distribution $F^{(1)}_{k,t}$ to drift over time. Assumption 4 ensures that neither nulls nor non-nulls become too rare over time. Assumption 5 guarantees the optimal score is identifiable.
Theorem 5 (Optimality of online score selection).
Suppose Assumptions 2-4 hold.
  1. (i)
    If $t+n\geq 1+1/(1-\rho)$, then for any $\varepsilon_{1}\in\left(0,\sqrt{\underline{\pi}/2}\right)$ and $\varepsilon_{2}\in(0,1-\rho)$,
    $\Pr\left(|{\mathcal{M}}^{\rm EWMA}(k,{\mathcal{D}}_{t})-{\mathbb{E}}[F^{k}_{0}(V({\bf X}_{t};k))\mid\theta_{t}=1]|\leq\varepsilon_{1}+\varepsilon_{2}+\frac{2\gamma}{(1-\overline{\pi})(1-\rho)}+\frac{2}{(t+n-1)\underline{\pi}}\right)$
    $\geq 1-3\exp\left\{-(t+n-1)\underline{\pi}\varepsilon_{1}^{2}\right\}-4\exp\left\{-\frac{\varepsilon_{2}^{2}({1-\overline{\pi}})^{2}}{6(1-\rho)}\right\}.$
    Further assume that Assumption 5 holds. Then the following results in (ii)-(iii) hold.
  2. (ii)
    If $t+n\geq 1+{16}/({c\underline{\pi}})$ and ${\gamma}/({1-\rho})\leq{(1-\overline{\pi})c}/{32}$, we have the finite-sample error bound for the score selection
    $\Pr(\hat{k}_{t}\neq k_{t}^{*}\mid\theta_{t}=1)\leq\;3K\exp\!\left\{-(t+n-1)\underline{\pi}\frac{c^{2}}{64}\right\}+4K\exp\!\left\{-\frac{c^{2}(1-\overline{\pi})^{2}}{384(1-\rho)}\right\}.$
  3. (iii)
    Moreover, the score selection consistency for all non-null points holds as
    $\Pr(\exists i\leq t:\theta_{i}=1,\ \hat{k}_{i}\neq k_{i}^{*})\leq\frac{3K(1-\underline{\pi})\exp\!\left\{-n\underline{\pi}{c^{2}}/{64}\right\}}{1-\exp\!\left\{-\underline{\pi}{c^{2}}/{64}\right\}}+4K(1-\underline{\pi})t\exp\!\left\{-\frac{c^{2}(1-\overline{\pi})^{2}}{384(1-\rho)}\right\}.$
    In particular, if $(1-\rho)\log(t)\to 0$, $\gamma/(1-\rho)\to 0$ and $n\to\infty$, $\Pr(\exists i\leq t:\theta_{i}=1,\ \hat{k}_{i}\neq k_{i}^{*})\to 0$.
Theorem 5 (i) shows that the EWMA criterion concentrates around the oracle power proxy. The bound decomposes naturally into: an estimation term scaling with $t+n$, a non-null effective size term in EWMA operation by $(1-\rho)^{-1}$ and a drift bias of order $\gamma/(1-\rho)$ reflecting how well the EWMA estimator can track a drifting non-null distribution. Parts (ii)–(iii) translate this concentration into optimal score selection guarantees. The rate conditions reflect the classic EWMA trade-off: $\rho$ must be sufficiently close to 1 to ensure an adequate effective sample size, yet small enough to allow the procedure to adapt to distribution drift.
5 Numerical Experiments
@@
In this section, we present extensive synthetic experiments to demonstrate the validity and efficiency of our proposed methods. First, Section 5.1 studies GAIF and Adaptive GAIF in the non-conformal setting. Second, Sections 5.2 and 5.3 present the results for OCTF and Opt-OCTF under the conformal setting, respectively. Following the setup in Robertson et al. [2023], we fix $\lambda=0.5$, $\gamma_{j}\propto j^{-1.6}$ for all $j\in\mathbb{N}$, and $s_{0}=\alpha/2$ throughout.
@@ 5.1 Synthetic experiments under non-conformal settings
We begin by presenting experiments on online multiple testing problems under three scenarios. Throughout, let $\pi_{1}$ denote the non-null proportion parameter in the data-generating model; its value varies across experiments.
  • Scenario I (Testing with Gaussian observations) We simulate $T$ independent test statistics $Z_{t}\sim N(\mu_{t},1)$ and test hypotheses $\mathbb{H}_{0t}:\mu_{t}=0$ using one-sided $p$-values $p_{t}=\Phi(-Z_{t})$, where $\Phi$ is the standard Gaussian CDF. The signal strengths $\mu_{t}$ follow a mixture model:
@@
    where the random variable $F_{1}\sim N(2.5,1)$.
  • Scenario II (Testing with Beta alternatives) We generate $p$-values according to the following model:
@@
  • Scenario III (Testing under local dependence) We simulate correlated test statistics $(Z_{1},\dots,Z_{T})^{\top}\sim N(\bm{\mu},\Sigma)$ where $\bm{\mu}=(\mu_{1},\ldots,\mu_{T})^{\top}$ with $\mu_{t}=2$ for randomly a fraction $\pi_{1}$ of indices and $\mu_{t}=0$ otherwise. The covariance matrix $\Sigma$ has a block-diagonal structure: coordinates are grouped into blocks of size $n_{\mathrm{block}}=10$, with within-block correlation $\rho=0.8$ and diagonal elements as 1. We test hypotheses $\mathbb{H}_{0t}:\mu_{t}=0$ using one-sided $p$-values $p_{t}=\Phi(-Z_{t})$.
We set the stopping time as $T=1000$ and the FDR level at $\alpha=0.1$. We evaluate the performance via empirical FDR and power averaged over 500 independent replications. Appendix E confirms that mFDR and FDR show similar trends.
@@
Firstly, we consider the full and instant feedback setting, and compare our methods SF and LF with state-of-the-art algorithms for online FDR control, namely LOND [Javanmard and Montanari, 2015], LORD++, and SAFFRON, using their default parameters ($\lambda=0.5$ chosen for SAFFRON). Figure 2 shows results for varying non-null proportion $\pi_{1}\in[0.1,0.8]$ under Scenarios I and II. All benchmarks ensure valid FDR control across different $\pi_{1}$ under both Scenarios. Our feedback-based SF and LF methods enhance detection power while maintaining valid online FDR control: SF significantly outperforms SAFFRON, while LF yields higher power than LORD++ and surpasses SAFFRON when $\pi_{1}$ is small. In contrast, both LORD++ and LOND remain conservative across different values of $\pi_{1}$. The plots have almost invisible error bars, as the variation across the 500 runs is negligible.
@@
Figure 2: Results for Scenario I and Scenario II under full and instant feedback: FDR and Power at stopping time with non-null proportion $\pi_{1}$ ranging from $0.1$ to $0.8$ across $500$ replications; The black dashed lines denote the FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
Figures 3 present results under a positive local dependence structure in Scenario III. The baseline methods—LOND, LORD++, and SAFFRON—are included for comparison. The proposed procedures, $\text{SF}_{\rm dep}$ and $\text{LF}_{\rm dep}$, correspond to $\text{Ada\text{-}GAIF}_{\rm dep}$ and $\rm GAIF_{\rm dep}$ under full and instant feedback, respectively. In addition, we include the existing methods $\text{SAFFRON}_{\rm dep}$ and $\text{LORD}_{\rm dep}$ from Zrnic et al. [2021]. We find that SF, LF, and SAFFRON fail to control the FDR under dependence. In contrast, the dependence-aware procedures $\mathrm{SF}_{\mathrm{dep}}$ and $\mathrm{LF}_{\mathrm{dep}}$ successfully achieve valid FDR control and attain substantially higher power than their feedback-ignoring counterparts, $\mathrm{SAFFRON}_{\mathrm{dep}}$ and $\mathrm{LORD}_{\mathrm{dep}}$.
Figure 3: Results for Scenario III under full and instant feedback: FDR and Power at stopping time with non-null proportion $\pi_{1}$ ranging from 0.1 to 0.8 across 500 replications. The black dashed line indicates the target FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
Next, we evaluate the performance of our proposed methods under other feedback settings: bandit and instant, and full and delayed. Taking the classical methods LOND, LORD++, and SAFFRON as benchmarks, we implement the proposed LF-BI and SF-BI under bandit and instant feedback setting. The results in Scenario I are shown in Figure 4. All methods maintain valid FDR control. Importantly, bandit feedback continues to enhance detection power, although the improvement is less pronounced than that under full and instant feedback.
Figure 4: Results for Scenario I under bandit and instant feedback: FDR and Power at stopping time with non-null proportion $\pi_{1}$ ranging from 0.1 to 0.8 across 500 replications. The black dashed line indicates the target FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
For the full and delayed feedback setting, we consider the delay $d\in\{10,100,200\}$ and compare the proposed SF-FD and LF-FD with the baselines without consideration of feedback. Figure 5 displays the results in Scenario I. As the delay $d$ increases, the detection power of our methods decreases slightly, yet they still outperform the baselines by a substantial margin while maintaining valid FDR control.
Figure 5: Results for Scenario I under full and delayed feedback: FDR and power at stopping time with non-null proportion $\pi_{1}$ ranging from 0.1 to 0.8 across 500 replications. The black dashed line indicates the target FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.
@@ 5.2 Results for online conformal testing with feedback
  • Scenario IV: Data is from $\mathbf{X}\mid Y=0\sim\mathcal{N}_{4}\left(\bm{\mu}_{1},\mathbf{I}_{4}\right)$, and $\mathbf{X}\mid Y=1\sim\mathcal{N}_{4}\left(\bm{\mu}_{2},\mathbf{I}_{4}\right)$, where $\bm{\mu}_{1}=(1,0,0,0)^{\top},\bm{\mu}_{2}=(0,0,-2,-2)^{\top}$. The target region is $\mathcal{A}=\{1\}$.
We set the stopping time to $T=1000$ and the target $\operatorname{FDR}$ level to $\alpha=0.2$. The initial calibration size is $n_{\text{cal}}=1000$. We consider the score as $V(\mathbf{X})=1-\hat{\mu}(\mathbf{X})$, where the predictive model $\hat{\mu}(\mathbf{X})$ is a neural network classifier trained on $n_{\text{tr}}=1000$. Appendix E reports additional results for varying $n_{\mathrm{cal}}$, alternative training algorithms, and a regression setting.
We evaluate the performance of the proposed LFS and SFS and implement the benchmarks LF, SF, LOND, LORD++ and SAFFRON within the OCTF workflow with their own testing levels $\{\alpha_{t}\}$. Figure 6 reports the empirical FDR and power for varying non-null proportions $\pi_{1}\in[0.1,0.8]$. All methods control the FDR below the nominal level $\alpha$. As expected, the feedback-enhanced procedures achieve clear power gains. It is noted that SF consistently achieves the highest power, while LF also performs competitively and attains higher power than SAFFRON. SFS and LFS outperform LORD++ and LOND owing to the incorporation of feedback information. This confirms that the safe conformal construction remains effective in practice, despite discarding certain rejections under the non-null for finite-sample guarantee.
Figure 6: Results for Scenario IV: FDR and power at stopping time with non-null proportion $\pi_{1}$ ranging from 0.1 to 0.8 across 500 replications. The black dashed lines denote the FDR level $\alpha=0.2$. Shaded areas show $\pm 1$ standard error.
@@ 5.3 Results for optimized online conformal testing with feedback
We now demonstrate the effectiveness of the proposed score selection strategy from Section 4.3. The goal here is to examine whether feedback can be used not only for valid testing, but also for adaptive score selection when non-null distributions shift.
We employ the binary classification example in Scenario IV with $\bm{\mu}_{1}=(2,0,0,0)^{\top},\bm{\mu}_{2}=(0,0,-2,-2)^{\top}$. For the historical labeled dataset, the non-null proportion is fixed as $0.2$. For the testing data, we consider a sine pattern shifts, where the non-null proportion varies as $\pi_{t}=\operatorname{Pr}(Y_{t}=1)=\{\sin(8\pi t/T)+1\}/4$, oscillating between 0 and 0.5.
@@
For the implementation of the score selection, we consider $K=3$ candidate models: random forest (RF), neural network (NN), and support vector machine (SVM) and employ the EWMA criterion ${\mathcal{M}}^{\text{EWMA}}_{t}(k,{\mathcal{D}}_{t})$ with exponential weight parameter $\rho=0.95$. To reduce computational cost, the auxiliary non-null $p$-value is computed by a sliding window of fixed length $L=200$: for $t>L$, we define ${\mathcal{C}}_{1t}=\{i:t-L\leq i<t,\theta_{i}=1\}$; for $t\leq L$, we use the full set ${\mathcal{C}}_{1t}=\{i<t:\theta_{i}=1\}$.
To evaluate the score-selection step, we compare the proposed optimized rule with random score selection under different testing procedures. Here, “Opt-” means that the score model is selected adaptively by the proposed EWMA criterion, while “Ran-” means that the score model is chosen uniformly at random from the candidate library. For example, Opt-LFS denotes the optimized version of LFS, whereas Ran-LFS uses the same testing rule with random score selection.
@@
Figure 7 shows the results under distribution shift. All methods successfully control empirical FDR around the nominal level $\alpha=0.1$. In terms of power, the optimized (Opt) variants consistently outperform their randomly selected (Ran) counterparts, implying that optimized score selection is useful under distribution shift. In addition, the LFS-based and SFS-based procedures exhibit lower power than LF-based and SF-based, respectively. However, in the next section, we find that the SF-based method may suffer from FDR inflation on real data due to the lack of finite-sample theoretical guarantees. Meanwhile, additional results for scenarios without distribution shifts are presented in Appendix E, and also highlight the effectiveness of our optimized score selection strategy.
@@
Figure 7: Results for Scenario V (sine pattern shifts): $\operatorname{FDR}(t)$ and $\operatorname{Power}(t)$ across different time $t$. The black dashed lines denote the FDR level $\alpha=0.1$. Shaded areas show $\pm 1$ standard error.